### v2

# Token Maturation: Autoregressive Language Generation via Continuous Token Dynamics

Abstract
Standard autoregressive language models collapse uncertainty at every generation step by committing to discrete tokens through immediate sampling. This premature discretization underlies well-known failure modes, including degenerate repetition loops in greedy decoding and a heavy reliance on heuristic sampling strategies.
We introduce Token Maturation, a continuous autoregressive framework in which tokens evolve as vector-valued trajectories prior to discretization. Rather than sampling from a categorical distribution at each step, the model resolves uncertainty through a deterministic dynamical process in embedding space, deferring discrete commitment until the representation has geometrically stabilized.
We show that this formulation mitigates degeneration intrinsically: Token Maturation generates coherent and diverse text under fully deterministic decoding (argmax), without repetition penalties, temperature scaling, or stochastic sampling. Moreover, we identify a novel convergence behavior in which token representations stabilize spatially while predictive entropy remains high, challenging the common assumption that commitment requires probability concentration. We propose continuous token dynamics with delayed commitment as an alternative formulation of autoregressive generation that exposes structural regularities obscured by immediate discretization.
autoregressive generation, continuous representations, delayed discretization, contrastive learning
1 Introduction
Figure 1: Immediate commitment vs. token maturation. (A) Standard decoding commits irreversibly at each step. (B) Token maturation maintains a continuous “liquid tail” (blue) that evolves geometrically, deferring commitment until convergence.
Autoregressive language models based on the Transformer architecture generate text by predicting a categorical distribution over tokens at each step (Vaswani et al., 2017). In modern implementations, this prediction is parameterized by a softmax over a fixed vocabulary, forcing the model to immediately commit to a discrete token via sampling or greedy selection. Once a token is selected, the decision becomes irreversible and fully conditions all subsequent generation.
While effective in practice, this design enforces early discretization of uncertainty. Continuous structure present in the model’s internal representations is collapsed into a categorical choice at every step, and uncertainty must be handled indirectly through token-level sampling heuristics. This coupling between prediction and commitment limits the ways in which uncertainty can be expressed and manipulated during generation.
Phenomenon: Template Attractors Prompt: "The meaning of life is..."
Run 1: "...eternal happiness," says Dr. James Wilson, director of Psychiatry...
Run 2: "...eternal peace," says Dr. R. Ehrlichman, director of Stanford Med...
Run 3: "...to know thyself," says Dr. David Hahn, director of Neuroscience...
Figure 2: Emergence of Template Attractors. Instead of collapsing to a single token sequence, Token Maturation converges to a semantic template (Quote $\rightarrow$ Attribution $\rightarrow$ Expert Title). The model maintains diversity in specific entities (names/places) while strictly adhering to the structural attractor.
In this work, we propose an alternative interface between prediction and commitment based on token maturation. Rather than committing to a discrete token at every generation step, we represent tokens as continuous vectors that evolve over time before discretization. Generation remains autoregressive and causal, but discretization is delayed: the model predicts trajectories in embedding space, allowing uncertainty to be represented geometrically and maintained throughout the maturation process until discrete commitment. A discrete token is committed only once the corresponding representation has sufficiently stabilized. Importantly, this final discretization step serves solely as an interface to the vocabulary and does not define the generative policy itself. Moreover, token maturation does not require a monotonic reduction in predictive entropy: discrete commitment can emerge even when entropy remains approximately constant throughout the maturation process.
This formulation yields a model that is fully autoregressive yet fundamentally distinct from both standard probabilistic decoding and diffusion-based text generation. Unlike conventional autoregressive models, which discretize uncertainty at every step, token maturation maintains continuous uncertainty until commitment becomes unavoidable. Unlike diffusion models, which operate on entire sequences via global denoising, token maturation is local, causal, and incremental. As a result, uncertainty is handled within continuous dynamics rather than being collapsed prematurely into a token-level sampling decision. Figure 1 contrasts immediate commitment with token maturation.
A direct consequence of delayed discretization is the emergence of additional degrees of freedom during generation. Because token representations remain continuous prior to commitment, the model admits interventions that are ill-defined in standard discrete autoregressive models. These include injecting noise into historical token representations, applying temporal smoothing or exponential moving averages over past states, and perturbing intermediate trajectories without altering committed tokens. Such interventions act on the continuous dynamics rather than on the discrete sampling process and provide structured ways to explore and stabilize generation trajectories without requiring entropy collapse. While not required for correct generation, they are naturally supported by the proposed framework and provide mechanisms for controlling stability and diversity that are unavailable in purely discrete models.
We study token maturation through a series of controlled experiments designed to isolate the effect of delayed discretization. We analyze the behavior of continuous token evolution under varying perturbation levels.
Contributions.
The main contributions of this work are:
  • We introduce token maturation, a continuous-variable, autoregressive language generation framework with delayed discretization.
  • We show that delayed discretization enables well-defined interventions on continuous token histories, such as noise injection and temporal smoothing, which are not naturally expressible in standard discrete autoregressive models.
  • We provide a mechanistic analysis of uncertainty resolution in continuous token space, identifying stable and collapse regimes under controlled perturbations.
  • We demonstrate a minimal instantiation using a standard Transformer architecture, confirming the feasibility of autoregressive language generation purely via continuous vector dynamics, without softmax-based decoding.
2 Related Work
2.1 Autoregressive Prediction and Continuous Relaxations
Standard autoregressive models commit to discrete tokens immediately via greedy or stochastic sampling (Vaswani et al., 2017; Radford et al., 2019; Holtzman et al., 2019). While methods like speculative decoding (Leviathan et al., 2023) optimize this process, they retain the immediate commitment assumption. Continuous relaxations (e.g., Gumbel-Softmax) (Jang et al., 2016; Maddison et al., 2016) approximate discrete choices for differentiability during training but still require discrete selection at inference. In contrast, Token Maturation treats tokens as evolving continuous vectors, deferring discretization until geometric convergence is achieved.
2.2 Diffusion and Hybrid Generative Models
Diffusion-based text generation typically relies on non-autoregressive, global denoising updates (Lou et al., 2023; Nie et al., 2025; Li et al., 2022), which differ fundamentally from our causal, left-to-right formulation. Recent hybrid architectures like Coconut (Hao et al., 2024) and TiDAR (Liu et al., 2025) introduce continuous latent states for reasoning or draft refinement. However, these models maintain the standard semantics of immediate token commitment at each step. Our approach uniquely decouples prediction from commitment, using continuous dynamics as the primary generative mechanism rather than as an auxiliary reasoning aid.
2.3 Contrastive Learning for Generation
Contrastive objectives (e.g., InfoNCE) are standard for representation learning (Chen et al., 2020; Gao et al., 2021) but rarely drive autoregressive generation. We repurpose contrastive learning to prevent regression-to-the-mean collapse in the absence of a categorical softmax likelihood. This anchors continuous predictions to semantic identities without forcing premature discretization.
3 Autoregressive token maturation
We introduce a continuous-variable formulation of autoregressive language modeling in which tokens are represented and generated as vectors in embedding space, and discrete commitment is deferred through a process we refer to as token maturation. This section formalizes the representation, generation dynamics, and training objective underlying the proposed framework.
3.1 Continuous Token Representation
Let $\mathcal{V}$ denote a discrete vocabulary of size $|\mathcal{V}|$, and let $E\in\mathbb{R}^{|\mathcal{V}|\times d}$ be a fixed embedding matrix, where each row $e_{i}\in\mathbb{R}^{d}$ corresponds to a token embedding. We assume embeddings are $\ell_{2}$-normalized and scaled to a fixed radius $R$.
Rather than predicting a categorical distribution over $\mathcal{V}$, the model predicts continuous vectors $z_{t}\in\mathbb{R}^{d}$ at each position $t$. Discrete tokens are recovered only at commitment time by projecting continuous vectors onto the vocabulary embedding set.
3.2 Autoregressive Vector Prediction
Given a sequence of previously committed token vectors $\{z_{1},\dots,z_{t-1}\}$, the model predicts a continuous vector $\hat{z}_{t}$ via an autoregressive function
$\hat{z}_{t}=f_{\theta}(z_{1},\dots,z_{t-1}),$ (1)
where $f_{\theta}$ is implemented as a causal Transformer operating directly in embedding space. Importantly, $\hat{z}_{t}$ is not immediately discretized and may evolve over time before commitment.
3.3 Conditioning on Maturation State
To enable the model to behave appropriately at different stages of the maturation process, we condition on two quantities: the noise level $\alpha$ at each position, and the tail length $K$.
Noise-level conditioning.
Figure 3: Architecture of the Continuous Autoregressive Backbone. Unlike standard LLMs, our model operates entirely in continuous space. Input vectors combine committed prefix tokens with noisy “liquid” tail tokens. The model is conditioned on the noise level $\alpha$ (via additive bias) and the tail length $K$ (via FiLM modulation), ensuring the network distinguishes between stable context and maturing predictions.
Each position in the sequence is associated with a noise level $\alpha_{t}\in[0,1]$, indicating how corrupted or uncertain the corresponding vector is. We embed $\alpha_{t}$ using a sinusoidal positional encoding (as in diffusion models) followed by a learned MLP, and add the result to the token representation:
$h_{t}\leftarrow h_{t}+\mathrm{MLP}(\mathrm{SinEmb}(\alpha_{t})).$ (2)
This allows the model to distinguish between committed tokens ($\alpha\approx 1$) and uncertain tail tokens ($\alpha\approx 0$).
Tail-length conditioning.
We further condition the model on the current tail length $K$ via feature-wise linear modulation (FiLM). A learned embedding of $K$ is projected to produce scale and shift parameters $(\gamma,\beta)$, which modulate the hidden representations:
$h\leftarrow(1+\gamma)\odot h+\beta.$ (3)
This global conditioning allows the model to adjust its predictions based on how much context is committed versus uncertain.
3.4 Token Maturation
To decouple prediction from commitment, we maintain a maturation buffer of length $K$, referred to as the liquid tail. At any generation step, the model maintains a sequence
$(z_{1},\dots,z_{t-K},\tilde{z}_{t-K+1},\dots,\tilde{z}_{t}),$
where the final $K$ vectors are uncommitted and continuously updated.
At each step, predicted vectors are iteratively refined according to
$\tilde{z}_{i}\leftarrow\tilde{z}_{i}+\alpha_{i}(\hat{z}_{i}-\tilde{z}_{i}),$ (4)
where $\alpha_{i}\in(0,1]$ controls the maturation rate. Earlier positions in the tail are updated more aggressively, while newly introduced vectors evolve slowly, resulting in gradual stabilization over time.
This process allows uncertainty to be expressed geometrically as distance in embedding space and resolved incrementally rather than through instantaneous sampling.
3.5 Discrete Commitment via Projection
Once a token vector reaches the front of the maturation buffer, it is committed by projection onto the embedding matrix:
$x_{t}=\arg\max_{i\in\mathcal{V}}\langle z_{t},e_{i}\rangle.$ (5)
The committed vector is then replaced by its corresponding embedding $e_{x_{t}}$ and becomes part of the fixed autoregressive context. Although commitment uses an argmax operation, stochasticity arises implicitly through the continuous maturation dynamics rather than explicit sampling.
3.6 Training Objective
Figure 4: Training with Simulated Maturation. To train the continuous dynamics without discrete ground truth at every step, we inject position-dependent noise into the tail vectors. The model is trained with a hybrid objective: MSE Loss (left) ensures geometric convergence to the target embedding, while InfoNCE Loss (right) utilizes negative sampling to maintain discriminative structure and prevent representation collapse.
A pure regression objective on continuous vectors leads to mode averaging and collapse. To stabilize training and align continuous predictions with discrete token identity, we combine a mean-squared error objective with a contrastive loss.
Given a predicted vector $\hat{z}_{t}$ and its ground-truth embedding $e_{x_{t}}$, we minimize
$\mathcal{L}_{\text{reg}}=\|\hat{z}_{t}-e_{x_{t}}\|_{2}^{2},$ (6)
alongside a contrastive InfoNCE loss
$\mathcal{L}_{\text{NCE}}=-\log\frac{\exp(\langle\hat{z}_{t},e_{x_{t}}\rangle/\tau)}{\sum_{j\in\mathcal{N}}\exp(\langle\hat{z}_{t},e_{j}\rangle/\tau)},$ (7)
where $\mathcal{N}$ is a set of negative samples and $\tau$ is a temperature parameter.
The final training objective is
$\mathcal{L}=\mathcal{L}_{\text{reg}}+\lambda\mathcal{L}_{\text{NCE}}.$ (8)
This contrastive component prevents collapse toward frequent tokens and anchors continuous predictions to discrete semantic identities without requiring a softmax likelihood.
4 Training and Generation
This section describes how the proposed model is trained and how autoregressive generation is performed at inference time. Although training and generation operate under different constraints, both are governed by the same underlying token maturation dynamics.
4.1 Training with Simulated Maturation
During training, the model is exposed to partially matured token representations to encourage robustness to uncertainty and to align training dynamics with inference-time behavior. Given a ground-truth token sequence $(x_{1},\dots,x_{T})$, we first map tokens to their embedding representations $(e_{x_{1}},\dots,e_{x_{T}})$.
To simulate the presence of a liquid tail, we perturb a suffix of length $K$ by mixing the ground-truth embeddings with isotropic noise in embedding space. Earlier tokens remain fixed, while later tokens are progressively corrupted, mimicking different stages of maturation. This procedure exposes the model to inputs ranging from fully committed tokens to highly uncertain representations.
The model is trained to predict the next-step continuous vector $\hat{z}_{t+1}$ given the current sequence of committed and uncommitted vectors, using the combined regression and contrastive objective described in Section 3.
Loss weighting.
To prevent the model from overweighting highly corrupted positions where the target is inherently ambiguous, we weight the loss at each position by $(1-\alpha_{t})$, where $\alpha_{t}$ is the noise level. Positions with low noise (near commitment) contribute more to the gradient, while highly uncertain positions contribute less.
4.2 Noise Injection and Stability
Noise injection during training serves two complementary purposes. First, it regularizes the model by preventing over-reliance on exact embedding vectors. Second, it approximates the distribution of uncommitted token states encountered during generation.
Importantly, noise is bounded and scaled such that vector norms evolve gradually over time. This ensures that uncertainty is resolved through maturation rather than abrupt stochastic jumps, and avoids the training–inference mismatch commonly encountered when noise is injected only at sampling time.
4.3 Autoregressive Generation
At inference time, generation proceeds autoregressively from left to right. Given an initial prompt, the corresponding token embeddings are inserted into the sequence as committed vectors. A liquid tail of length $K$ is initialized with low-norm random vectors, representing highly uncertain token states.
At each generation step, the model predicts updated continuous vectors for the entire sequence. Vectors in the liquid tail are updated according to the maturation rule, while committed tokens remain fixed. Once a vector reaches the front of the liquid tail, it is discretized via projection onto the vocabulary embedding matrix and committed permanently.
This process yields a stream of discrete tokens, while internally maintaining a continuous representation that evolves over time.
Classifier-free guidance.
At inference time, we optionally apply classifier-free guidance (CFG) to sharpen predictions toward the conditioned context. We compute two forward passes: a conditional pass using the full causal mask, and an unconditional pass using a tail-only mask that prevents tail tokens from attending to history. The final prediction is a weighted combination:
$\hat{z}_{t}=\hat{z}_{t}^{\text{uncond}}+s\cdot(\hat{z}_{t}^{\text{cond}}-\hat{z}_{t}^{\text{uncond}}),$ (9)
where $s\geq 1$ is the guidance scale. This encourages generated tokens to be more consistent with the committed context.
4.4 Generation Algorithm
Algorithm 1 summarizes the proposed autoregressive generation procedure.
Algorithm 1 Autoregressive Generation with Token Maturation
0: Prompt embeddings $(e_{x_{1}},\dots,e_{x_{n}})$, tail length $K$, guidance scale $s$
1: Initialize committed sequence $\mathbf{z}_{1:n}\leftarrow(e_{x_{1}},\dots,e_{x_{n}})$
2: Initialize liquid tail $\tilde{\mathbf{z}}_{n+1:n+K}$ with random low-norm vectors
3: Construct alpha profile: $\alpha_{t}=1$ for $t\leq n$, fading from $\alpha_{\max}$ to 0 over tail
4:while not end-of-sequence do
5:  $\hat{\mathbf{z}}^{\text{cond}}\leftarrow f_{\theta}(\mathbf{z},\boldsymbol{\alpha},K)$ {full causal mask}
6:  $\hat{\mathbf{z}}^{\text{uncond}}\leftarrow f_{\theta}(\mathbf{z},\boldsymbol{\alpha},K)$ {tail-only mask}
7:  $\hat{\mathbf{z}}\leftarrow\hat{\mathbf{z}}^{\text{uncond}}+s\cdot(\hat{\mathbf{z}}^{\text{cond}}-\hat{\mathbf{z}}^{\text{uncond}})$ {CFG}
8:  Update tail: $\tilde{z}_{i}\leftarrow\tilde{z}_{i}+\alpha_{i}(\hat{z}_{i}-\tilde{z}_{i})$ for $i$ in tail
9:  Commit front token: $x_{n+1}\leftarrow\arg\max_{j}\langle\tilde{z}_{n+1},e_{j}\rangle$
10:  Replace: $z_{n+1}\leftarrow e_{x_{n+1}}$, append new embryo to tail
11:  $n\leftarrow n+1$, update $\boldsymbol{\alpha}$
12:end while
13:return Generated token sequence $(x_{1},\dots,x_{n})$
4.5 Computational Considerations
The proposed framework introduces minimal overhead relative to standard autoregressive generation. The primary additional cost arises from maintaining and updating the liquid tail, which scales linearly with the tail length $K$. In practice, we find small values of $K$ sufficient to capture maturation dynamics, keeping inference costs comparable to conventional decoding methods.
5 Experiments
We evaluate token maturation through experiments designed to validate its core claims: that coherent text can be generated without entropy collapse, that tail length controls diversity, and that learned embeddings adapt to the continuous prediction task.
5.1 Experimental Setup
We train a 24-layer causal Transformer with hidden dimension 1024 and 16 attention heads, operating directly in embedding space. The model is trained on the FineWeb 10BT dataset (Penedo et al., 2024) for 600K steps with batch size 8 and gradient accumulation over 4 steps, yielding an effective batch size of 32. Token embeddings are initialized from GPT-2 and optionally fine-tuned during training. We use the combined MSE + InfoNCE objective described in Section 3, with 256 negative samples and a logit scale of 20.
Unless otherwise specified, we use a liquid tail of length $K=16$ during generation. All experiments use deterministic decoding (argmax) without temperature scaling or nucleus sampling.
5.2 Comparison with Standard Greedy Decoding
A well-documented failure mode of autoregressive language models is that greedy (argmax) decoding produces degenerate, repetitive text (Holtzman et al., 2019). In practice, coherent generation requires sampling-based methods such as temperature scaling, top-k, or nucleus sampling, or explicit penalties that discourage repetition.
We evaluate whether Token Maturation resolves this failure mode intrinsically. We compare three configurations: (1) GPT-2 with greedy decoding, (2) GPT-2 with greedy decoding and repetition penalty of 1.1, and (3) Token Maturation with greedy decoding. All models generate up to 80 tokens from 20 diverse prompts. We measure distinct-n (ratio of unique n-grams) and rep-n (fraction of repeated n-grams).
Table 1 shows the results. GPT-2 with pure greedy decoding collapses almost immediately: distinct-2 drops to 0.30 and 90% of generations contain trigram loops. Adding a repetition penalty restores diversity (distinct-2 = 0.999), but this requires an external intervention that explicitly penalizes repeated tokens during decoding.
Token Maturation achieves perfect diversity (distinct-2 = 1.0, rep-2 = 0.0) using pure argmax decoding, without any penalties or sampling. This suggests that the continuous maturation process intrinsically prevents the degeneration that plagues standard autoregressive decoding. The model does not need external correction because uncertainty is resolved geometrically in embedding space rather than being forced through a premature discrete commitment.
Table 1: Repetition metrics under greedy decoding. GPT-2 requires repetition penalty to avoid degeneration; Token Maturation achieves comparable quality intrinsically. Results averaged over 20 prompts, 80 tokens max.
Token Mat. GPT-2 GPT-2 + Rep.
Metric (greedy) (greedy) (penalty=1.1)
Distinct-1 $\uparrow$ 0.974 0.228 0.936
Distinct-2 $\uparrow$ 1.000 0.301 0.999
Rep-2 $\downarrow$ 0.000 0.699 0.001
Rep-3 $\downarrow$ 0.000 0.665 0.000
Trigram loop (%) $\downarrow$ 0 90 0
We emphasize that repetition penalties are effective practical tools; our claim is not that they are unnecessary, but that they compensate for a failure mode induced by immediate discretization.
5.3 Coherent Generation without Entropy Collapse
Figure 5: Geometric Stabilization vs. Discrete Jumps. Visualization of the maturation trajectory in embedding space. While standard decoding (inset) forces immediate discrete jumps, our model allows the token representation to traverse a continuous manifold. The trajectory stabilizes geometrically (convergence to a specific region) before the final commitment, allowing the model to resolve semantic ambiguity smoothly over time.
A central prediction of our framework is that discrete commitment can occur without a corresponding reduction in predictive entropy. To test this, we track the entropy of the model’s implicit distribution over vocabulary tokens throughout the maturation process.
At each maturation step, we compute the cosine similarity between the current vector and all vocabulary embeddings, apply a softmax with temperature $\tau=1$, and measure the resulting entropy.
Figure 6 shows entropy trajectories for representative generation runs. Contrary to the expectation that commitment requires certainty, we observe that entropy remains approximately constant ($H\approx 3.9$ nats) throughout maturation, decreasing only marginally before the final snap. Despite this sustained uncertainty, generated text is syntactically coherent and topically consistent.
This finding supports our central claim: the model converges not to a single token, but to a region in embedding space where multiple semantically appropriate tokens reside at similar distances. Commitment emerges from geometric proximity rather than probability concentration.
Figure 6: Left: Entropy throughout token maturation for four representative tokens. Despite progressing through 24 maturation steps, entropy remains constant at approximately 3.9 nats—the model never collapses to certainty before commitment. Right: Top candidate for token “Dr” at each step, showing exploration through semantically diverse alternatives despite constant entropy. This demonstrates that commitment emerges from geometric convergence, not probability concentration.
5.4 Tail Length Controls Diversity
The liquid tail length $K$ determines how many maturation steps each token undergoes before commitment. We hypothesize that shorter tails preserve more of the initial randomness, yielding diverse outputs, while longer tails allow convergence toward a deterministic trajectory.
To test this, we generate 50 continuations from the same prompt under varying tail lengths ($K\in\{1,4,16,64\}$), using identical model weights but different random initializations for the tail.
5.5 Emergent Template Structure
A striking qualitative observation is that Token Maturation often converges to a structural template before committing to specific lexical content. Figure 2 illustrates this phenomenon: given the prompt “The meaning of life is”, the model consistently generates text following the pattern:
“[Philosophical claim],” says Dr. [Name], director of [Institution]…
Across multiple runs with different random initializations, the high-level structure remains stable while the specific details—the doctor’s name, institutional affiliation, and quoted content—vary freely.
This behavior is not naturally expressible under immediate token commitment, where structure must emerge incrementally through irreversible lexical decisions. Conventional decoding commits to each token immediately and irrevocably; the model cannot “plan” a structural template because it has no mechanism to defer lexical commitment while exploring structural alternatives. Each token choice constrains all future choices, forcing structure to emerge incrementally rather than being selected holistically.
In contrast, Token Maturation’s continuous dynamics naturally support this separation. During the maturation process, token representations evolve in embedding space before discretization. We hypothesize that the model first converges to regions corresponding to structural or syntactic roles (e.g., “a quoted statement” followed by “attribution to an expert”), and only later resolves the specific tokens that instantiate this structure. The constant-entropy observation from Section LABEL:sec:entropy supports this interpretation: the model maintains uncertainty over multiple valid instantiations of the same structural template, rather than collapsing to a single deterministic sequence.
This emergent planning behavior—structure before content—is a direct consequence of decoupling prediction from commitment. It suggests that continuous token dynamics may offer a path toward language generation with more global coherence, without requiring explicit planning modules or latent structure variables.
5.6 Qualitative Examples
Figure 7 shows representative generation samples with the liquid tail visualized. blue tokens indicate pre-commitment states. We observe that tokens often remain ambiguous up to the commitment stage.
Figure 7: Generation interface showing token maturation in action. Left: Committed text (white) followed by the liquid tail (cyan) containing uncommitted tokens that will mature over subsequent steps. Bottom: Live metrics including entropy ($H=3.91$), confirming sustained uncertainty. Right: Top candidates for the next commitment, showing near-uniform scores over semantically appropriate alternatives (psychology, professor, psychiatrist, neuro). Despite high entropy, all candidates are contextually relevant—the model converges to a semantic region rather than a single token.
6 Discussion and Conclusion
We introduced Token Maturation, a framework that reformulates autoregressive generation by explicitly decoupling prediction from commitment. Standard language models collapse uncertainty into a discrete decision at every step; in contrast, our model maintains a continuous token state that stabilizes geometrically over time. This design enables uncertainty to be represented as a trajectory in embedding space rather than a categorical distribution, yielding smooth semantic drift under ambiguity.
Geometric Argmax and the Absence of Repetition.
Our results challenge the common intuition that argmax decoding is inherently greedy or repetitive. We demonstrated that when argmax is applied after stochastic evolution in continuous space, it behaves analogously to a Gumbel-max process, producing diverse and coherent text. Crucially, Token Maturation exhibits no degenerate repetition loops—a pervasive failure mode in standard LLMs (Holtzman et al., 2019). We hypothesize that the continuous dynamics naturally prevent convergence to repetitive attractors, suggesting that repetition in standard models arises partly from the premature nature of discrete commitment.
Dynamics and Stability.
Unlike diffusion-based approaches, Token Maturation remains strictly autoregressive and causal, preserving the incremental nature of generation. A key practical finding is the necessity of contrastive objectives (such as InfoNCE) to prevent representation collapse. Regression alone tends to produce overly smooth outputs; learning a meaningful continuous manifold requires explicit pressure to preserve discriminative structure alongside geometric convergence.
Future Directions.
While this work focuses on conceptual clarity, it opens a broad design space for temporally extended generation. Future work should explore scaling these dynamics to high-capacity models, implementing adaptive maturation schedules that allocate more computation to ambiguous tokens, and integrating symbolic constraints. We view delayed commitment not merely as a decoding trick, but as a fundamental reinterpretation of the autoregressive process.
Appendix A Appendix
A.1 Embedding Geometry Adapts to Continuous Prediction
When embeddings are fine-tuned during training, we observe systematic reorganization of the embedding space. Figure 8 visualizes the drift between frozen GPT-2 embeddings and learned embeddings after 600K training steps.
Several patterns emerge:
  • Stable tokens: Years (1978, 1987, 1992) and common function words exhibit minimal drift, suggesting GPT-2’s geometry is already suitable for these tokens.
  • High-drift tokens: Punctuation, Unicode symbols, and rare code fragments drift substantially, indicating that the original embeddings poorly served continuous prediction for these tokens.
  • Semantic reorganization: Nearest-neighbor relationships shift in interpretable ways. For instance, “Python” moves from proximity to other programming languages (Java, PHP) toward proximity to programming culture tokens (Lisp, Emacs, Unix).
This reorganization occurs without explicit supervision on embedding structure, emerging purely from the continuous prediction objective.
Figure 8: Embedding drift from frozen GPT-2 to learned embeddings. Left: distribution of drift across vocabulary. Right: tokens with highest drift are predominantly punctuation and rare symbols.
A.2 Classifier-Free Guidance Reveals Interpretable Lookahead
Classifier-free guidance (CFG) interpolates between conditional and unconditional predictions, typically used to sharpen generation toward the prompt. In our framework, CFG has an additional effect: it pulls uncommitted tail vectors toward the manifold of coherent text, making intermediate states semantically interpretable.
Figure 9 compares tail states with and without CFG ($s=1$ vs. $s=2$). Without guidance, tail tokens project to seemingly random vocabulary items with no coherent relationship to the context or to each other. With guidance, tail tokens form interpretable sequences that reflect forward planning: topics, syntactic continuations, and semantic themes become visible before commitment.
This suggests that CFG does not merely sharpen final predictions, but actively shapes the geometry of the maturation trajectory. The liquid tail becomes a window into the model’s latent “reasoning”—a form of interpretable lookahead that is unavailable in standard autoregressive generation.
Figure 9: Effect of CFG on tail interpretability. Without CFG (top), tail tokens appear as incoherent noise. With CFG (bottom), tail tokens form semantically meaningful lookahead, revealing the model’s implicit forward planning.